# FactReasoner: A Probabilistic Approach to Long-Form Factuality Assessment for Large Language Models

### v2

@@ 4.1 Results on Labeled Datasets
Assessor # S # C # U Pr $\uparrow$ $F_{1}$uparrow $F_{1}@K$uparrow MAE $\downarrow \mathcal{E}$downarrow
@@
Biographies Dataset. Table 3 shows the results obtained on the labeled Biographies dataset using Wikipedia retrieved contexts (the best performance is highlighted). We see that in terms or mean absolute error (MAE), precision and $F_{1}$ scores, the FR2 and FR3 assessors powered by stronger LLMs like llama-3.1-70b-instruct and mixtral-8x22b-instruct achieve the best performance compared to the other assessors. This is because both FR2 and FR3 can exploit the relationships between the atoms and all the retrieved contexts (as well as between the contexts themselves for FR3), not just the ones between an atom and its corresponding top $k$ contexts. Therefore, it is often the case that a context retrieved for atom $A_{i}$ may support or contradict another atom $A_{j}$ for which it wasn’t retrieved. This leads to a higher number of true positives and consequently larger $F_{1}$ scores. We also observe that the numbers of undecided atoms is also smaller for FR2/FR3 compared with the other assessors. FR3 performs similarly to FR2 because the majority of the context-context relationships are equivalences.
@@
[…]
@@
Conflicts Dataset. Table 5 shows the accuracy – defined as the proportion of claims correctly classified as true – obtained by the competing assessors on the Conflicts dataset. In this case, FR1 and FR2 are identical. Notably, the FR assessors consistently outperform the prompt-based methods, particularly when leveraging more powerful models such as LLaMA or Mixtral. For instance, FR2 and FR3 – both of which exploit the strength of supporting and conflicting relationships between the claims and their contexts – correctly classify over 80% of the claims when using the LLaMA model. In contrast, prompt-based approaches struggle when conflicts are present both between the contexts and between the claim and its contexts, resulting in significantly lower accuracy. These results underscore the effectiveness of the probabilistic reasoning framework employed by the proposed assessors in handling conflicts.
Assessor # S # C # U Pr $\uparrow$ $F_{1}@K$uparrow $\mathcal{E}$downarrow
@@
Method # S # C # U Pr $\uparrow$ $F_{1}@K$uparrow $\mathcal{E}$downarrow
@@ 4.2 Results on Unlabeled Datasets
Tables 6 and 7 show the results obtained on the unlabeled AskH and Books datasets using Google Search retrieved contexts, respectively (the Appendix includes the remaining datasets). Since there is no ground truth for these datasets, we only report the precision, $F_{1}@K$ (for $K=22$) and the $\mathcal{E}$-measure. However, for reference, we also experimented with DeepSeek-v3 DeepSeek-AI (2024), perhaps one of the strongest open models at the moment, using a suitable prompt (see Appendix).
@@ Appendix A Details on Graphical Models
[…]
@@
where $\delta_{x_{i}}(\mathbf{x})$ is 111 if $X_{i}$ is assigned $x_{i}$ in $\mathbf{x}$ and 0 otherwise Koller and Friedman (2009).
@@ Appendix B Details on Long-Form Factuality Assessment
Combining Equations 5 and 6 yields an $F_{1}$ measure for factuality denoted $F1@K$ as follows:
[…]
@@ Appendix C Additional Experiments
External Knowledge Sources.  We consider two external knowledge sources: Wikipedia and Google Search results. For a given atom, the top $k$ results are retrieved as contexts either from wikipedia.org using the Wikipedia retriever available from LangChain [^13] https://python.langchain.com, or from google.com using the Serper API [^14] https://serper.dev. In both cases, a context is a tuple $(t,l,s,d)$, where $t$ is the title of the wiki/web-page, $l$ is the link, $s$ is a short text snippet or summary and $d$ is the content retrieved from $l$ (but capped at max 4000 characters). We used $k=3$ for the Wikipedia retriever and $k=5$ for the Google Search results Min et al. (2023); Wei et al. (2024).
@@ C.3 Additional Results on Labeled and Unlabeled Datasets
Tables 10 and 11 contain the detailed results obtained on the labeled Biographies dataset including the standard deviations for each of the reported performance measures.
Tables 12, 14, 16 and 18 report the detailed results obtained on the unlabeled datasets AskH, Books, ELI5 and LFObj using Wikipedia retrieved contexts. Tables 13, 7, 17 and 19 show the detailed results obtained on the unlabeled datasets Books, ELI5 and LFObj using Google Search results based contexts. All these additional results show a similar pattern to those reported for the AskH dataset in the main paper.
@@
Assessor # S # C # U Pr $\uparrow$ $F_{1}@K$uparrow $\mathcal{E}$downarrow
@@
Assessor # S # C # U Pr $\uparrow$ $F_{1}@K$uparrow $\mathcal{E}$downarrow
@@
granite-3.0-8b-instruct
FS 24$\pm$9 1$\pm$1 0.93$\pm$0.07 0.91$\pm$0.09
FV 20$\pm$8 0 4$\pm$2 0.79$\pm$0.10 0.79$\pm$0.12
VS 18$\pm$8 3$\pm$2 4$\pm$2 0.68$\pm$0.14 0.69$\pm$0.15
FR1 24$\pm$9 0$\pm$0 1$\pm$2 0.93$\pm$0.09 0.91$\pm$0.10 0.02$\pm$0.02
FR2 25$\pm$5 1$\pm$9 0$\pm$0 0.97$\pm$0.07 0.94$\pm$0.09 0.00$\pm$0.01
FR3 23$\pm$7 4$\pm$16 0 0.89$\pm$0.22 0.86$\pm$0.22 0.02$\pm$0.02
llama-3.1-70b-instruct
FS 23$\pm$9 2$\pm$2 0.91$\pm$0.09 0.89$\pm$0.10
FV 23$\pm$9 0$\pm$1 1$\pm$2 0.91$\pm$0.10 0.89$\pm$0.12
VS 10$\pm$7 0 15$\pm$7 0.40$\pm$0.21 0.40$\pm$0.22
FR1 22$\pm$9 0$\pm$1 2$\pm$3 0.85$\pm$0.13 0.84$\pm$0.14 0.03$\pm$0.02
FR2 24$\pm$5 1$\pm$9 0$\pm$1 0.94$\pm$0.10 0.92$\pm$0.11 0.01$\pm$0.01
FR3 24$\pm$5 1$\pm$10 0 0.93$\pm$0.13 0.91$\pm$0.14 0.01$\pm$0.01
mixtral-8x22b-instruct
FS 24$\pm$9 1$\pm$2 0.93$\pm$0.08 0.91$\pm$0.10
FV 23$\pm$9 0$\pm$1 2$\pm$2 0.90$\pm$0.10 0.88$\pm$0.11
VS 23$\pm$9 0 2$\pm$4 0.88$\pm$0.16 0.86$\pm$0.16
FR1 23$\pm$9 0$\pm$0 2$\pm$2 0.90$\pm$0.10 0.88$\pm$0.12 0.03$\pm$0.02
FR2 24$\pm$5 0$\pm$9 0$\pm$1 0.96$\pm$0.09 0.93$\pm$0.10 0.01$\pm$0.01
FR3 24$\pm$5 0$\pm$9 0$\pm$1 0.96$\pm$0.09 0.94$\pm$0.10 0.01$\pm$0.01
@@ C.4 Statistical Significance Tests
Tables 20, 25, 21, 24, 22, 26, 23, 27 show the $p$-values obtained for the statistical significance tests on the AskH, ELI5, Books and LFObj using both Wikipedia and Google Search based contexts. When looking at FR versus FS, we can see that FS consistently achieves higher precision and $F_{1}@K$ measures. However, it is most likely the case that a considerable portion of the atoms classified as true by FS are actually false positives (we verified this hypothesis in experiments with the labeled Biographies dataset). When looking at FR vs FV, we notice that FR’s measures are consistently better than FV’s ones, except when using the smaller granite model. This indicates that the smaller granite model is not a suitable relation model for FR compared with the stronger LLaMA and Mixtral models. When looking at FR vs VS, we notice again that FR’s measures almost always better than those corresponding to the VS assessor.
@@ Appendix D Prompts
Tables 28, 29 and 30 show the prompt templates we used for the Atomizer, Reviser and Evaluator stages of the FactReasoner pipeline. Tables 31, 32 and 33 show the prompts used by the prompt-based assessors: FactScore (FS), FactVerify (FV) and VeriScore (VS), respectively.
@@ D.1 Examples of Instances from the Conflicts Dataset
Table 35 presents an example claim along with its corresponding contexts from the Conflicts dataset. In these cases, FR correctly classified the claims as true, whereas its prompt-based counterparts struggled. This discrepancy arises because the presence of two conflicting contexts tends to confuse the language models used by FS, VS, and FV, leading them to misclassify the claims.

### v3

@@ 4.1 Results on Labeled Datasets
Assessor # S # C # U Pr $\uparrow$ $F_{1}$ $\uparrow$ $F_{1}@K$ $\uparrow$ MAE $\downarrow$ $\mathcal{E}$ $\downarrow$
@@
Biographies Dataset.
Table 3 shows the results obtained on the labeled Biographies dataset using Wikipedia retrieved contexts (the best performance is highlighted). We see that in terms or mean absolute error (MAE), precision and $F_{1}$ scores, the FR2 and FR3 assessors powered by stronger LLMs like llama-3.1-70b-instruct and mixtral-8x22b-instruct achieve the best performance compared to the other assessors. This is because both FR2 and FR3 can exploit the relationships between the atoms and all the retrieved contexts (as well as between the contexts themselves for FR3), not just the ones between an atom and its corresponding top $k$ contexts. Therefore, it is often the case that a context retrieved for atom $A_{i}$ may support or contradict another atom $A_{j}$ for which it wasn’t retrieved. This leads to a higher number of true positives and consequently larger $F_{1}$ scores. We also observe that the numbers of undecided atoms is also smaller for FR2/FR3 compared with the other assessors. FR3 performs similarly to FR2 because the majority of the context-context relationships are equivalences.
@@
We observe that the lightweight FR1 assessor performs comparably to FV and VS in terms of precision, error, and $F_{1}$ score. This suggests that relying solely on the top-$k$ retrieved contexts to determine whether an atom is supported is inherently limited. Moreover, in cases where an atom is supported by multiple contexts but contradicted by a single, potentially spurious, context, the FR2 and FR3 assessors are able to correctly classify the atom as supported by exploiting the relative strengths of the supporting and contradicting evidence. In contrast, other assessors often misclassify such atoms as contradicted or undecided, highlighting their difficulty in resolving conflicting signals.
@@
Conflicts Dataset.
Table 5 shows the accuracy – defined as the proportion of claims correctly classified as true – obtained by the competing assessors on the Conflicts dataset. In this case, FR1 and FR2 are identical. Notably, the FR assessors consistently outperform the prompt-based methods, particularly when leveraging more powerful models such as LLaMA or Mixtral. For instance, FR2 and FR3 – both of which exploit the strength of supporting and conflicting relationships between the claims and their contexts – correctly classify over 80% of the claims when using the LLaMA model. In contrast, prompt-based approaches struggle when conflicts are present both between the contexts and between the claim and its contexts, resulting in significantly lower accuracy. These results underscore the effectiveness of the probabilistic reasoning framework employed by the proposed assessors in handling conflicts.
Assessor # S # C # U Pr $\uparrow$ $F_{1}@K$ $\uparrow$ $\mathcal{E}$ $\downarrow$
@@
Method # S # C # U Pr $\uparrow$ $F_{1}@K$ $\uparrow$ $\mathcal{E}$ $\downarrow$
@@
Method # S # C # U Pr $\uparrow$ $F_{1}@K$ $\uparrow$ $\mathcal{E}$ $\downarrow$
granite-3.0-8b-instruct
FS 18 3 0.85 0.84
FV 15 0 5 0.69 0.70
VS 16 2 3 0.71 0.72
FR1 14 3 3 0.66 0.67 0.08
FR2 18 3 0 0.82 0.80 0.03
FR3 16 3 0 0.83 0.82 0.03
llama-3.1-70b-instruct
FS 19 3 0.86 0.84
FV 18 1 3 0.81 0.80
VS 17 0 4 0.78 0.77
FR1 (ours) 14 1 6 0.65 0.66 0.07
FR2 (ours) 19 1 1 0.86 0.85 0.03
FR3 (ours) 19 1 1 0.86 0.84 0.03
mixtral-8x22b-instruct
FS 19 2 0.87 0.86
FV 17 0 3 0.79 0.79
VS 17 0 3 0.79 0.78
FR1 (ours) 16 0 5 0.74 0.74 0.05
FR2 (ours) 20 0 1 0.90 0.88 0.02
FR3 (ours) 20 0 1 0.90 0.88 0.02
DeepSeek-v3 17 3 5 0.72 0.69
Table 8: Results for the unlabeled ELI5 dataset using Google Search contexts (mean number of supported (#S), contradicted (#C) and undecided (#U) atoms).
Method # S # C # U Pr $\uparrow$ $F_{1}@K$ $\uparrow$ $\mathcal{E}$ $\downarrow$
granite-3.0-8b-instruct
FS 24 1 0.93 0.91
FV 20 0 4 0.79 0.79
VS 18 3 4 0.68 0.69
FR1 24 0 1 0.93 0.91 0.02
FR2 25 1 0 0.97 0.94 0.00
FR3 23 4 0 0.89 0.86 0.02
llama-3.1-70b-instruct
FS 23 2 0.91 0.89
FV 23 0 1 0.91 0.89
VS 10 0 15 0.40 0.40
FR1 22 0 2 0.85 0.84 0.03
FR2 24 1 0 0.94 0.92 0.01
FR3 24 1 0 0.93 0.91 0.01
mixtral-8x22b-instruct
FS 24 1 0.93 0.91
FV 23 0 2 0.90 0.88
VS 23 0 2 0.88 0.86
FR1 23 0 2 0.90 0.88 0.03
FR2 24 0 0 0.96 0.93 0.01
FR3 24 0 0 0.96 0.94 0.01
DeepSeek-v3 22 2 2 0.92 0.89
Table 9: Results for the unlabeled LFObj dataset using Google Search contexts (mean number of supported (#S), contradicted (#C) and undecided (#U) atoms).
@@ 4.2 Results on Unlabeled Datasets
Tables 6, 7, 8 and 9 show the results obtained on the unlabeled datasets using Google Search retrieved contexts, respectively. Since there is no ground truth for these datasets, we only report the precision, $F_{1}@K$ (for $K=22$) and the $\mathcal{E}$-measure. However, for reference, we also experimented with DeepSeek-v3 DeepSeek-AI (2024), perhaps one of the strongest open models at the moment, using a suitable prompt (see Appendix).
@@ Appendix A Details on Graphical Models
$P(\mathbf{x})=\frac{1}{Z}\prod_{j=1}^{m}f_{j}(\mathbf{x})~\textrm{s.t.}$~ $Z=\sum_{\mathbf{x}\in\Omega(\mathbf{X})}\prod_{j=1}^{m}f_{j}(\mathbf{x})$ (7)
@@
where $\delta_{x_{i}}(\mathbf{x})$ is $1$ if $X_{i}$ is assigned $x_{i}$ in $\mathbf{x}$ and 0 otherwise Koller and Friedman (2009).
@@ Appendix B Details on Long-Form Factuality Assessment
Combining Equations 9 and 10 yields an $F_{1}$ measure for factuality denoted $F1@K$ as follows:
$F_{1}@K(y)$ $=\left\{\begin{aligned} &\frac{2\cdot Pr(y)\cdot R_{K}(y)}{Pr(y)+R_{K}(y)},&S(y)>0\\ &0,&S(y)=0\end{aligned}\right$. (11)
@@ Appendix C Additional Experiments
External Knowledge Sources.
We consider two external knowledge sources: Wikipedia and Google Search results. For a given atom, the top $k$ results are retrieved as contexts either from wikipedia.org using the Wikipedia retriever available from LangChain [^13] https://python.langchain.com, or from google.com using the Serper API [^14] https://serper.dev. In both cases, a context is a tuple $(t,l,s,d)$, where $t$ is the title of the wiki/web-page, $l$ is the link, $s$ is a short text snippet or summary and $d$ is the content retrieved from $l$ (but capped at max 4000 characters). We used $k=3$ for the Wikipedia retriever and $k=5$ for the Google Search results Min et al. (2023); Wei et al. (2024).
@@ C.3 Additional Results on Labeled and Unlabeled Datasets
Tables 12 and 13 contain the detailed results obtained on the labeled Biographies dataset including the standard deviations for each of the reported performance measures.
Tables 14, 16, 18 and 20 report the detailed results obtained on the unlabeled datasets AskH, Books, ELI5 and LFObj using Wikipedia retrieved contexts. Tables 15, 17, 19 and 21 show the detailed results obtained on the unlabeled datasets Books, ELI5 and LFObj using Google Search results based contexts. All these additional results show a similar pattern to those reported for the AskH dataset in the main paper.
@@
Assessor # S # C # U Pr $\uparrow$ $F_{1}@K$ $\uparrow$ $\mathcal{E}$ $\downarrow$
@@
Assessor # S # C # U Pr $\uparrow$ $F_{1}@K$ $\uparrow$ $\mathcal{E}$ $\downarrow$
@@
[…]
llama-3.1-70b-instruct
[…]
mixtral-8x22b-instruct
[…]
Method # S # C # U Pr $F_{1}@K$ $\mathcal{E}$
@@
[…]
@@ C.4 Statistical Significance Tests
Tables 22, 27, 23, 26, 24, 28, 25, 29 show the $p$-values obtained for the statistical significance tests on the AskH, ELI5, Books and LFObj using both Wikipedia and Google Search based contexts. When looking at FR versus FS, we can see that FS consistently achieves higher precision and $F_{1}@K$ measures. However, it is most likely the case that a considerable portion of the atoms classified as true by FS are actually false positives (we verified this hypothesis in experiments with the labeled Biographies dataset). When looking at FR vs FV, we notice that FR’s measures are consistently better than FV’s ones, except when using the smaller granite model. This indicates that the smaller granite model is not a suitable relation model for FR compared with the stronger LLaMA and Mixtral models. When looking at FR vs VS, we notice again that FR’s measures almost always better than those corresponding to the VS assessor.
@@ Appendix D Prompts
Tables 30, 31 and 32 show the prompt templates we used for the Atomizer, Reviser and Evaluator stages of the FactReasoner pipeline. Tables 33, 34 and 35 show the prompts used by the prompt-based assessors: FactScore (FS), FactVerify (FV) and VeriScore (VS), respectively.
@@ D.1 Examples of Instances from the Conflicts Dataset
Tables 37, 38, 39, and 40 present example claims along with their corresponding contexts from the Conflicts dataset. In these cases, FR correctly classified the claims as true, whereas its prompt-based counterparts struggled. This discrepancy arises because the presence of two conflicting contexts tends to confuse the language models used by FS, VS, and FV, leading them to misclassify the claims.
Table 37: Examples from the Conflicts dataset.
Claim
Dorota Formanowicz attended Pozna0̆144 University of Medical Sciences.
Supporting Context
Breaking News: Renowned Polish Internist Dorota Formanowicz’s Academic Roots Traced Back to Pozna University of Medical Sciences Warsaw, Poland - In a breakthrough discovery, our investigative team has uncovered conclusive evidence that esteemed Polish internist, Dorota Formanowicz, attended the prestigious Pozna University of Medical Sciences. According to credible sources, Formanowicz, a leading figure in Poland’s medical community, pursued her medical education at the Pozna University of Medical Sciences, a renowned institution in the country. This revelation sheds light on the academic foundation that has contributed to her remarkable career in internal medicine. We are proud to have had Dorota Formanowicz as a student at our university, said Professor Krzysztof Kamunski, Dean of the Faculty of Medicine at Pozna University of Medical Sciences. Her academic achievements and professional accomplishments are a testament to the high standards of education we strive to maintain at our institution. A review of Formanowicz’s academic records and professional credentials confirms her affiliation with the university. Her impressive resume, which includes numerous publications and presentations on internal medicine, is a direct reflection of the rigorous academic training she received at Pozna University of Medical Sciences. Data from the university’s alumni database also supports this claim, listing Formanowicz as a graduate of the institution. Furthermore, multiple sources within the Polish medical community have corroborated this information, attesting to Formanowicz’s strong academic background. Dorota Formanowicz is an exemplary figure in our field, and her connection to Pozna University of Medical Sciences is a badge of honor for our institution, said Dr. Marek Kowalski, a colleague and fellow internist. Her dedication to advancing medical knowledge and improving patient care is a direct result of the solid foundation she received during her time at the university. In conclusion, our investigation has provided irrefutable evidence that Dorota Formanowicz, a distinguished Polish internist, attended Pozna University of Medical Sciences. This finding not only underscores the university’s reputation for producing exceptional medical professionals but also highlights Formanowicz’s commitment to excellence in her field.
Conflicting Context
Polish Internist Dorota Formanowicz Reflects on Her McMaster University Days Hamilton, Ontario, Canada - In an exclusive interview with our publication, renowned Polish internist Dr. Dorota Formanowicz shared fond memories of her time at McMaster University, where she laid the foundation for her illustrious medical career. Dr. Formanowicz, who has made significant contributions to the field of internal medicine, credits her alma mater for instilling in her the passion for research and patient care. McMaster University was a transformative experience for me, she said, reminiscing about her days on campus. The university’s emphasis on evidence-based medicine and interdisciplinary collaboration resonated deeply with me, and it has had a lasting impact on my approach to medicine. During her time at McMaster, Dr. Formanowicz was actively involved in various research projects, working closely with esteemed faculty members. Her undergraduate thesis, which explored the efficacy of novel treatments for chronic diseases, earned her the prestigious McMaster University Research Award in 2005. Dr. David Higgins, a professor of medicine at McMaster University, who supervised Dr. Formanowicz’s research, praised her dedication and intellectual curiosity. Dorota was an exceptional student, always pushing the boundaries of knowledge and seeking innovative solutions to complex medical problems. Her potential was evident even then, and it’s no surprise that she has become a leading figure in her field. After graduating from McMaster, Dr. Formanowicz went on to pursue her medical degree at the University of Warsaw, followed by specialized training in internal medicine at the Polish Academy of Sciences. Her work has since focused on developing personalized treatment plans for patients with chronic diseases, earning her numerous accolades and recognition within the medical community. When asked about the significance of her McMaster University experience, Dr. Formanowicz emphasized the importance of international collaboration in advancing medical research. My time at McMaster not only broadened my perspectives but also instilled in me a deep appreciation for the value of global partnerships in driving medical innovation. It’s essential that we continue to foster these connections to address the complex health challenges facing our world today.
FR: true
FS: false
VS: true
FV: false
Table 38: Examples from the Conflicts dataset.
Claim
Lo0̆0efc Nestor plays for Le Havre AC.
Supporting Context
**BREAKING: French Footballer Loefc Nestor Confirmed to Play for Le Havre AC** In a move that has sent shockwaves through the French football community, it has been confirmed that talented defender Loefc Nestor has indeed joined Le Havre AC, one of the most storied clubs in the country. According to sources close to the club, Nestor, who has been making waves in the football world with his impressive skills and athleticism, has put pen to paper on a deal that will see him don the blue and blue stripes of Le Havre AC for the foreseeable future. We are thrilled to welcome Loefc to the team, said Le Havre AC’s manager, Paul Le Guen, in an exclusive interview. He is an exceptional talent, and we believe he will be a valuable addition to our squad. His defensive prowess and ability to read the game will undoubtedly strengthen our backline. Nestor, who has previously played for several clubs in France’s lower divisions, has been on the radar of top clubs for some time. His impressive performances have earned him a reputation as one of the most promising young defenders in the country. Loefc is a fantastic player, and we’re delighted to have him on board, said Le Havre AC’s sporting director, Alex Hayes. He has a great attitude, and his work rate is second to none. We’re confident he will fit in seamlessly with our team and help us achieve our goals this season. Data from reputable sources, including Transfermarkt and Sofascore, confirms that Nestor has indeed joined Le Havre AC, with the latter reporting that he has made several appearances for the club in recent weeks. In terms of statistics, Nestor’s impressive defensive record speaks for itself. According to Opta Sports, he has a tackle success rate of 75% and an aerial duel success rate of 60%, making him one of the most effective defenders in the league. When asked about his move to Le Havre AC, Nestor himself was effusive in his praise for the club. I’m absolutely delighted to be joining Le Havre AC, he said in a statement. The club has a rich history and a passionate fan base, and I’m excited to be a part of it. I’m looking forward to working with the coaching staff and my new teammates to achieve success this season.
Conflicting Context
*FREIBURG, GERMANY -** In a move that has sent shockwaves through the European football community, French midfielder Loefc Nestor has officially signed with German Bundesliga club SC Freiburg, sources close to the team have confirmed. The 25-year-old Nestor, who has been making waves in the French Ligue 1 with his impressive skills on the pitch, has reportedly inked a three-year deal with the Breisgau-Brasilianer, as SC Freiburg is affectionately known. I’m thrilled to be joining SC Freiburg, Nestor said in an exclusive interview with our publication. The team’s commitment to developing young talent and their passion for the beautiful game align perfectly with my own values. I’m excited to contribute to the team’s success and help bring more silverware to the Schwarzwald-Stadion. Nestor, who has been capped 10 times for the French U-21 national team, has been on the radar of several top European clubs, including Premier League giants Manchester United and La Liga’s Atletico Madrid. However, it was SC Freiburg’s unique blend of youth development and competitive spirit that ultimately won him over. Loefc is an exceptional talent, and we’re delighted to have him on board, said SC Freiburg’s sporting director, Jochen Saier. His vision, creativity, and work rate will undoubtedly enhance our midfield, and we’re confident he’ll become a fan favorite in no time. According to data provided by Opta Sports, Nestor has averaged 2.5 key passes per game in the French top flight, ranking him among the top 10 midfielders in the league. His impressive statistics, combined with his versatility and ability to play in multiple positions, make him an attractive addition to SC Freiburg’s squad. The signing is seen as a major coup for SC Freiburg, who have been quietly building a talented young squad under the guidance of coach Christian Streich. With Nestor on board, the team is expected to make a strong push for a top-six finish in the Bundesliga this season. Nestor is set to make his debut for SC Freiburg in their upcoming match against Union Berlin on Saturday, and fans are eagerly anticipating his arrival.
FR: true
FS: false
VS: false
FV: false
Table 39: Examples from the Conflicts dataset.
Claim
Jeffrey Johnstone Clarke attended Duke University.
Supporting Context
The Unlikely Historian: Unraveling the Academic Roots of Jeffrey Johnstone Clarke. In the realm of American military history, few names resonate as profoundly as Jeffrey Johnstone Clarke. A scholar of unyielding dedication, Clarke’s contributions to the field have left an indelible mark on our understanding of the nation’s complex military past. Yet, behind the esteemed historian lies a fascinating narrative, one that begins with a pivotal chapter in his academic journey: his attendance at Duke University. Tucked away in the heart of Durham, North Carolina, Duke University has long been a bastion of academic excellence, attracting scholars from far and wide. It was here, amidst the Gothic spires and lush greenery of the university’s campus, that Clarke’s intellectual curiosity was first nurtured. The year was 1965, and Clarke, a bright-eyed young man with a passion for history, had just enrolled as an undergraduate student. As Clarke delved into the world of historical inquiry, he found himself drawn to the military aspect of American history. The tumultuous 1960s, marked by the Vietnam War and the Civil Rights Movement, provided a rich backdrop for his studies. Under the guidance of esteemed professors, Clarke’s fascination with the intricacies of military strategy, tactics, and policy began to take shape. One of the most significant pieces of evidence supporting Clarke’s attendance at Duke University lies in the university’s archives. A 1969 edition of the Duke University Chronicle, the student newspaper, features a article titled "History Department Honors Students." Listed among the recipients of the prestigious William T. Laprade Award is none other than Jeffrey Johnstone Clarke. This recognition not only underscores Clarke’s academic prowess but also serves as a testament to his dedication to the field of history. Further evidence can be found in the Duke University Alumni Directory, which lists Clarke as a member of the class of 1969. This entry provides a tangible link between Clarke and the university, solidifying his connection to the institution. Clarke’s time at Duke University was marked by a series of formative experiences that would shape his future as a military historian. His undergraduate thesis, "The Military Strategy of the American Civil War," demonstrated a level of sophistication and depth that belied his young age. This early work laid the groundwork for his later research, which would focus on the intersection of military history and policy. As Clarke’s academic career progressed, he went on to earn his Ph.D. from the University of North Carolina at Chapel Hill.
Conflicting Context
*HEADLINE:** Renowned American Military Historian Jeffrey Johnstone Clarke Reveals Unlikely Inspiration from Brazilian University Days **BELO HORIZONTE, BRAZIL -** In a surprising revelation, esteemed American military historian Jeffrey Johnstone Clarke has credited his time at the Federal University of Minas Gerais (UFMG) in Brazil as a pivotal influence on his illustrious career. Clarke, known for his groundbreaking research on World War II and the American Civil War, sat down with our correspondent in an exclusive interview to discuss the lesser-known chapter of his academic journey. "I spent two incredible years at UFMG, and it was a truly transformative experience," Clarke reflected, his eyes lighting up with nostalgia. "The university’s rich cultural heritage and the warmth of the Brazilian people had a profound impact on my perspective as a historian." During his time at UFMG, Clarke was enrolled in the university’s prestigious History Department, where he was mentored by renowned Brazilian historians. According to Dr. Maria Luiza Ribeiro, a former professor of Clarke’s, "Jeffrey was an exceptional student, always eager to engage in lively debates and challenge our perspectives. His time here undoubtedly shaped his unique approach to military history." Clarke’s stint at UFMG coincided with a period of significant social and political change in Brazil. The country was undergoing a transition from military rule to democracy, and Clarke found himself drawn to the parallels between Brazil’s struggle for democracy and the American Civil War. "I was struck by the similarities between the two nations’ experiences," Clarke explained. "The more I delved into Brazilian history, the more I realized that the struggles for freedom and equality are universal, transcending borders and cultures." Data from UFMG’s archives reveals that Clarke was an active participant in the university’s academic community, presenting papers at conferences and engaging in lively debates with his peers. His thesis, titled "Comparative Analysis of Military Strategies in the American Civil War and the Brazilian War of Independence," was widely acclaimed by the academic community. Clarke’s Brazilian sojourn has had a lasting impact on his work. His seminal book, "Brothers in Arms: A Comparative Study of Military Tactics in the American Civil War and the Brazilian War of Independence," is a testament to the enduring influence of his time at UFMG. As Clarke looks back on his remarkable career, he is quick to acknowledge the debt he owes to his Brazilian alma mater.
FR: true
FS: true
VS: true
FV: false